A 2-C charge experiences a force of 20N when put at a certain location in space. The electric field at that location is
1
Expert's answer
2013-02-11T08:45:48-0500
The formula for the& experienced& force is

F=E*q

where E is the electric field. Hence, electric field is

E= F/q = 20N/2C& = 10 N/C = 10 V/m